Could I have the answer for this question please.
7-3(5t-10)=67
is the answer t=6 or t=2
when calculating do I use -3x15 or 3x15. Could you give me some algebra tips for this. Thanks

Answer by AnlytcPhil(1116)

Could I have the answer for this question please. 
7-3(5t-10)=67 
is the answer t=6 or t=2
when calculating do I use -3x15 or 3x15. Could you give me some algebra tips
for this. Thanks 

The first one!  When removing parentheses by the distributive rule,
always take the sign in front of the multiplier along with the multiplier.

7 - 3(5t-10) = 67

Write down the 7

7

Then taking the sign with the 3, multiply -3 by 5t and get -15t.  Write that
down, so you have

7 - 15t

Now multiply the -3 by the -10, getting +30, so write than next. Now you have

7 - 15t + 30

Then copy " = 67 "

7 - 15t + 30 = 67 

Combine the 7 with the 30 and get 37

    37 - 15t = 67

Subtract 37 from both sides and get

        -15t = 30

Then divide both sides by -15

       -15t     30
      覧覧覧 = 覧覧
       -15     -15

           t = -2

Edwin
AnlytcPhil@aol.com